OpenStudy (anonymous):

What is the distance between the points (-2, -3) and (1, 1)? -3 5 7 1

OpenStudy (anonymous):

5

OpenStudy (anonymous):

dis= sqrt((x1-x2)2- (y1-y2)2) =sqrt((-2_1)2+(-3-1)2) =sqrt(6+16) =st(25) =5